# An extension of May’s Theorem to three alternatives: axiomatizing Minimax voting

### v2

@@ 3 Main result
Block preservation says that adding a block of voters who submit one copy of each linear order does not cause any alternative that would have won before to lose after. For three-alternative profiles, this simply means that adding six voters with the rankings $abc$, $acb$, $bac$, $bca$, $cab$, and $cba$ does not cause any previous winner to lose. Minimax satisfies a stronger block invariance axiom according to which the set of winners cannot change at all, but again it suffices for our main result to assume only the weaker axiom. [^9] We do not know of other papers discussing block preservation or block invariance, but they can be derived from other well-known axioms, such as the conjunction of anonymity, neutrality, and reinforcement: if $X(\mathbf{P})=X(\mathbf{P}^{\prime})$ and $V(\mathbf{P})\cap V(\mathbf{P}^{\prime})=\varnothing$, then $F(\mathbf{P})\cap F(\mathbf{P}^{\prime})\neq\varnothing$ implies $F(\mathbf{P}+\mathbf{P}^{\prime})=F(\mathbf{P})\cap F(\mathbf{P}^{\prime})$. Anonymity and neutrality together imply that if $\mathbf{P}^{\prime}$ consists of exactly one copy of each linear order of $X(\mathbf{P})$, then $F(\mathbf{P}^{\prime})=X(\mathbf{P}^{\prime})$ (as does the cancellation property of Young 1974), in which case reinforcement yields $F(\mathbf{P}+\mathbf{P}^{\prime})=F(\mathbf{P})\cap X(\mathbf{P}^{\prime})=F(\mathbf{P})$. However, no voting method that always picks the Condorcet winner uniquely when one exists can satisfy reinforcement (see Zwicker 2016, Prop. 2.5, and cf. Remark 4.26 of Holliday and Pacuit 2023a for doubts about the plausibility of reinforcement), whereas block invariance is widely satisfied.
@@
As far as we know, the only voting method in the literature that violates homogeneity is the Dodgson method (see Brandt 2009). [^10] Brandt (2009) gives a profile $\mathbf{P}$ for which $Dodgson(\mathbf{P})\neq Dodgson(3\mathbf{P})$; observing that $Dodgson(2\mathbf{P})\not\subseteq Dodgson(4\mathbf{P})$ yields a violation of (inclusion) homogeneity in our formulation. Violations of block preservation are slightly more common, coming not only from Dodgson but also from Bucklin (see Hoag and Hallett 1926), Coombs (Coombs 1964, Grofman and Feld 2004), and Mediancentre-Borda (Cervone et al. 2012). [^11] For Bucklin, let $\mathbf{P}$ contain $1$ voter with the ranking $abc$ and $3$ with $cab$. Then since $c$ has a majority of first-place votes, $c$ is the Bucklin winner. However, where $\mathbf{P}^{\prime}$ consists of each of the six linear orders over $\{a,b,c\}$, in $\mathbf{P}+\mathbf{P}^{\prime}$ no alternative receives a majority of first-place votes, but $a$ and $c$ both receive a majority of first-or-second-place votes, and $a$ has more first-or-second-place votes ($8$) than $c$ does ($7$), so $a$ is the Bucklin winner. For Coombs, let $\mathbf{P}$ contain $2$ voters with the ranking $abc$, $2$ with $cba$, $2$ with $cab$, and $1$ with $bac$. Then $c$ has a majority of first-place votes, so $c$ is the Coombs winner. But where $\mathbf{P}^{\prime}$ consists of each of the six linear orders over $\{a,b,c\}$, in $\mathbf{P}+\mathbf{P}^{\prime}$ no alternative receives a majority of first-place votes, and $c$ has the most last-place votes, so $c$ is eliminated in the first round (and then $a$ wins). Finally, for Mediancentre-Borda, start with the profile labeled “Irresolute OWM Failure” in the simulation accompanying Cervone et al. 2012 at https://omega.math.union.edu/research/2010-05-voting/; then adding one copy of each linear order causes the original winner to lose. However, clearly all scoring rules (Young 1975) and voting methods based on majority margins (see, e.g., Holliday and Pacuit 2023a) satisfy block preservation. Moreover, the methods cited above that violate block preservation also violate immunity to spoilers, so block preservation is not doing any special work by ruling out those methods.
We are now ready to state and prove our main result.
@@
Now suppose $\mathbf{P}$ is a three-alternative profile with no Condorcet winner. The margin graph of a profile $\mathbf{P}$ is the weighted directed graph with an edge from $a$ to $b$ if $a$ has a positive margin over $b$, weighted by that margin. It follows that the margin graph of $\mathbf{P}$ is of one of the following four forms, where the dashed arrows might not be present, due to a possibly zero margin:
@@
Note that here ‘$a$’, ‘$b$’, and ‘$c$’ above are variables—so in the ascending case, we simply call the alternative with the largest loss ‘$a$’, etc. To see that the above four cases are exhaustive, assuming there is no Condorcet winner, observe that if the number of edges in the margin graph is zero, then we are in symmetric case; if the number of edges is one, then we are in the ascending case with $0=n=m$; if the number of edges is two, then we are in either the Condorcet loser case or the ascending case or descending case with $0=n<m$; and if the number of edges is three, then we are in either the ascending, descending, or symmetric case.
@@
Case 3: the margin graph of $\mathbf{P}$ is as in the Condorcet loser case. In this case, $Minimax(\mathbf{P})=\{b,c\}$, and the argument that $F(\mathbf{P})\subseteq\{b,c\}$ is the same as in the first paragraph of the ascending case, only $\mathrm{Margin}_{\mathbf{P}}(a,b)$ might not be positive, which is irrelevant to the argument.
@@
In Section 2, we assumed in the tradition of Arrow (1963) that the domain of a voting method is the set of profiles of strict weak orders. However, inspection of the proofs of Facts 3.1, 3.3, 3.4, and 3.7 and of Theorem 3.8 shows that for these results, we do not actually need to assume any properties of a voter’s binary relation $\mathbf{P}(i)$. [^12] For an arbitrary binary relation $\mathbf{P}(i)$, when we say that $i$ “ranks $a$ uniquely first”, we simply mean that $(a,b)\in\mathbf{P}(i)$ for all $b\in X(\mathbf{P})\setminus\{a\}$ (and similarly for “ranks $a$ uniquely last”), not that $\mathbf{P}(i)$ forms a transitive ranking. In particular, all of these results also apply to the setting of linear ballots, where ties are not allowed, to the setting of truncated ballots, where some alternatives may be unranked (regarded as tied for last), and to the setting where instead of asking voters for a ranking, we simply ask for their pairwise choices (Do you prefer $a$ or $b$? Do you prefer $a$ or $c$? Do you prefer $b$ or $c$?), which allows for intransitivities in $\mathbf{P}(i)$. This latter point is especially important given the recent proposal of Foley and Maskin (2024) that in top-three general elections, we should ask voters to answer three pairwise comparison questions rather than to provide a ranking of the three alternatives. Theorem 3.8 continues to apply in this setting.
@@
Theorem 3.8 bears on a debate about how Minimax should be formulated. When an alternative $b$ has a positive margin over $a$, should we measure the size of $a$’s loss by the margin of $b$ over $a$, as we have, or rather by the support for $b$ over $a$, defined as the absolute number of voters who rank $b$ above $a$ (cf. Schulze 2011, p. 269)? If voters’ rankings are all strict linear orders, there is no difference, but in profiles allowing ties (even if only at the bottom of voters’ rankings), the two approaches are not equivalent. In fact, Minimax formulated in terms of support rather than margin violates positive involvement. Consider the following profile (where the number above each ranking indicates the number of voters with that ranking, and $b,c$ indicates a tie) with its margin graph (center) and support graph (right):
@@
Support-based Minimax says that $c$ has the “smallest loss”, so $c$ wins. But then as in Case 1a of the proof of Theorem 3.8, adding $3$ voters with the ranking $cba$ makes $b$ the Condorcet winner and hence the support-based Minimax winner, violating positive involvement. We take this to be an argument in favor of the margin-based formulation of Minimax instead of the support-based formulation.
Finally, we note that there are indeed proper refinements of Minimax satisfying the axioms in Theorem 3.8. Since the issue of how to break Minimax ties is much less important than constraining a voting method to choose from among the Minimax winners, we defer discussion of such refinements to the Appendix. However, it is noteworthy that with the refinement defined in the Appendix, we satisfy not only weak positive responsiveness but even full positive responsiveness as in Remark 2.2.
@@ 4 Independence of axioms
Let us now turn to the question of the independence of the axioms in Theorem 3.8. First, we claim that we can drop May’s axiom of anonymity from our list of axioms. Anonymity is only used in the proof of Theorem 3.8 to help derive that the voting method reduces to majority voting in two-alternative with a unique majority winner. But we can derive this result without anonymity as follows.
@@
Suppose $\mathbf{P}$ is a two-alternative profile with a unique majority winner. Hence there are $a,b\in X(\mathbf{P})$ such that $n$ voters rank $b$ above $a$, $n+k$ (for $k>0$) rank $a$ above $b$, and the remaining $\ell$ voters rank them as tied.
First suppose that $k>1$. Consider the profile $\mathbf{P}_{0}$ consisting of just $k-2$ of the voters from $\mathbf{P}$ who rank $a$ above $b$, still ranking $a$ above $b$, and all of the $\ell$ voters from $\mathbf{P}$ who ranked them as tied, still ranking them as tied. We claim that $F(\mathbf{P}_{0})=\{a\}$. For if $b\in F(\mathbf{P}_{0})$, then where $\mathbf{Q}_{0}$ is obtained from $\mathbf{P}_{0}$ by the $k-2$ voters who ranked $a$ above $b$ switching to rank $b$ above $a$, weak positive responsiveness applied $k-2$ times implies $F(\mathbf{Q}_{0})=\{b\}$, while neutrality implies $a\in F(\mathbf{Q}_{0})$, a contradiction. Hence $F(\mathbf{P}_{0})=\{a\}$. Now observe that there are exactly $n+2$ voters from $\mathbf{P}$ who ranked $a$ above $b$ who are not in $\mathbf{P}_{0}$. Let $\mathbf{P}_{1}$ by obtained from $\mathbf{P}_{0}$ by adding all of those $n+2$ voters, where $n+1$ of them still rank $a$ above $b$ but one of them—say, voter $i_{1}$—now ranks $b$ above $a$, and adding all $n$ of the original voters who ranked $b$ above $a$, still ranking $b$ above $a$. Then $\mathbf{P}_{1}$ is obtained from $\mathbf{P}_{0}$ by adding $n+1$ pairs of the rankings $ab$ and $ba$, so by block preservation, $F(\mathbf{P}_{0})=\{a\}$ implies $a\in F(\mathbf{P}_{1})$. Finally, observe that $\mathbf{P}$ can be obtained from $\mathbf{P}_{1}$ by voter $i_{1}$ moving $a$ from last place to first place, so $a\in F(\mathbf{P}_{1})$ implies $F(\mathbf{P})=\{a\}$ by weak positive responsiveness, in agreement with majority voting. For $k>2$, the relations between $\mathbf{P}$, $\mathbf{P}_{0}$, and $\mathbf{P}_{1}$ are shown in Figure 1.
Now suppose $k=1$. By homogeneity, $F(\mathbf{P})\subseteq F(2\mathbf{P})$. Then by the argument of the previous paragraph applied to $2\mathbf{P}$, for which the analogous $k$ value is $2$, we have $F(2\mathbf{P})=\{a\}$. Hence $F(\mathbf{P})=\{a\}$ in agreement with majority voting.∎
@@
$\mathbf{P}$ $h_{1}$dots $h_{n}$ $i_{1}$dots $i_{n+k}$ $j_{1}$dots $j_{\ell}$ $b$dots $b$ $a$dots $a$ $a,b$dots $a,b$ $a$dots $a$ $b$dots $b$
$\mathbf{P}_{0}$ $i_{n+3}$dots $i_{n+k}$ $j_{1}$dots $j_{\ell}$ $a$dots $a$ $a,b$dots $a,b$ $b$dots $b$
$\mathbf{P}_{1}$ $h_{1}$dots $h_{n}$ $\boldsymbol{i_{1}}$ $i_{2}$dots $i_{n+2}$ $i_{n+3}$dots $i_{n+k}$ $j_{1}$dots $j_{\ell}$ $b$dots $b$ $\boldsymbol{b}$ $a$dots $a$ $a$dots $a$ $a,b$dots $a,b$ $a$dots $a$ $\boldsymbol{a}$ $b$dots $b$ $b$dots $b$ $\scriptstyle n+1$ voters $\scriptstyle n+1$ voters voters from $\scriptstyle\mathbf{P}_{0}$
@@
All of the other axioms in Theorem 3.8 are independent of the other axioms in the following sense.
@@
For each of the axioms $A$ in Theorem 3.8 besides anonymity, there is a voting method $F$ on the domain of profiles with two or three alternatives such that $F$ satisfies all of the axioms in Theorem 3.8 except for $A$, and there is a profile $\mathbf{P}$ such that $F(\mathbf{P})\neq Minimax(\mathbf{P})$.
@@
Where $A$ is near immunity to spoilers, the Borda voting method clearly satisfies all the axioms except for near immunity to spoilers, which it violates in three-alternative profiles (see Holliday and Pacuit 2021a, Example 4.15); it is also easy to construct profiles with $Borda(\mathbf{P})\neq Minimax(\mathbf{P})$. ∎
@@
In $\mathbf{P}$, there is no Condorcet winner, so $\mbox{{CP}}=\{c\}$ since $c$ has the maximum plurality score of $5$. Then $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter with the ranking $cba$, but in $\mathbf{P^{\prime}}$, $b$ is the unique weak Condorcet winner, so $\mbox{{CP}}(\mathbf{P})=\{b\}$, violating positive involvement.
@@
Case 2: there is no weak Condorcet winner in $\mathbf{P}$. Then $a\in\mbox{{CP}}(\mathbf{P})$ implies that $a$ is among the alternatives with maximal plurality score in $\mathbf{P}$. Then clearly $a$ has the unique maximum plurality score in $\mathbf{P}^{\prime}$, and if moving from $\mathbf{P}$ to $\mathbf{P}^{\prime}$ creates a weak Condorcet winner, it must be $a$ alone, so in either case we have $\mbox{{CP}}(\mathbf{P})=\{a\}$.
@@
Observe that in $2\mathbf{P}$, the difference between the best and second best Minimax scores is 2, so $F(2\mathbf{P})=\{a\}$, violating homogeneity. Yet clearly $F$ satisfies anonymity, neutrality, block preservation (since adding a block of all linear orders does not change margins), and immunity to spoilers (since $F$ satisfies Condorcet consistency). It remains to show that $F$ satisfies weak positive responsiveness and positive involvement.
For positive involvement, for an arbitrary profile $\mathbf{P}$, suppose $x\in F(\mathbf{P})$ and that $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $x$ first. If $x\in Minimax(\mathbf{P})$, then $x\in Minimax(\mathbf{P}^{\prime})$ by positive involvement for Minimax, so $x\in F(\mathbf{P}^{\prime})$ by the definition of $F$. On the other hand, if conditions (I)-(III) hold and $x$ is the alternative with the second best Minimax score in $\mathbf{P}$, then since $x$ beats the Minimax winner head-to-head, in $\mathbf{P}^{\prime}$ the Minimax winner’s Minimax score increases by 1, while $x$’s Minimax score decreases by $1$, so $x\in Minimax(\mathbf{P}^{\prime})\subseteq F(\mathbf{P}^{\prime})$.

### v4

@@ 3 Main result
Block preservation says that adding a block of voters who submit one copy of each linear order does not cause any alternative that would have won before to lose after. For three-alternative profiles, this simply means that adding six voters with the rankings $abc$, $acb$, $bac$, $bca$, $cab$, and $cba$ does not cause any previous winner to lose. Minimax satisfies a stronger block invariance axiom according to which the set of winners cannot change at all, but again it suffices for our main result to assume only the weaker axiom. [^9] We do not know of other papers discussing block preservation or block invariance, but they can be derived from other well-known axioms, such as the conjunction of anonymity, neutrality, and reinforcement: if $X(\mathbf{P})=X(\mathbf{P}^{\prime})$ and $V(\mathbf{P})\cap V(\mathbf{P}^{\prime})=\varnothing$, then $F(\mathbf{P})\cap F(\mathbf{P}^{\prime})\neq\varnothing$ implies $F(\mathbf{P}+\mathbf{P}^{\prime})=F(\mathbf{P})\cap F(\mathbf{P}^{\prime})$. Anonymity and neutrality together imply that if $\mathbf{P}^{\prime}$ consists of exactly one copy of each linear order on $X(\mathbf{P})$, then $F(\mathbf{P}^{\prime})=X(\mathbf{P}^{\prime})$ (as does the cancellation property of Young 1974), in which case reinforcement yields $F(\mathbf{P}+\mathbf{P}^{\prime})=F(\mathbf{P})\cap X(\mathbf{P}^{\prime})=F(\mathbf{P})$. However, no voting method that always picks the Condorcet winner uniquely when one exists can satisfy reinforcement (see Zwicker 2016, Prop. 2.5, and cf. Remark 4.26 of Holliday and Pacuit 2023a for doubts about the plausibility of reinforcement), whereas block invariance is widely satisfied.
@@
As far as we know, the only voting method in the literature that violates homogeneity is the Dodgson method (see Brandt 2009). [^10] Brandt (2009) gives a profile $\mathbf{P}$ for which $Dodgson(\mathbf{P})\neq Dodgson(3\mathbf{P})$; observing in addition that $Dodgson(2\mathbf{P})\not\subseteq Dodgson(4\mathbf{P})$ yields a violation of (inclusion) homogeneity in our formulation. Violations of block preservation are slightly more common, coming not only from Dodgson but also from Bucklin (see Hoag and Hallett 1926), Coombs (Coombs 1964, Grofman and Feld 2004), and Mediancentre-Borda (Cervone et al. 2012). [^11] For Bucklin, let $\mathbf{P}$ contain $1$ voter with the ranking $abc$ and $3$ with $cab$. Then since $c$ has a majority of first-place votes, $c$ is the Bucklin winner. However, where $\mathbf{P}^{\prime}$ consists of each of the six linear orders over $\{a,b,c\}$, in $\mathbf{P}+\mathbf{P}^{\prime}$ no alternative receives a majority of first-place votes, but $a$ and $c$ both receive a majority of first-or-second-place votes, and $a$ has more first-or-second-place votes ($8$) than $c$ does ($7$), so $a$ is the Bucklin winner. For Coombs, let $\mathbf{P}$ contain $2$ voters with the ranking $abc$, $2$ with $cba$, $2$ with $cab$, and $1$ with $bac$. Then $c$ has a majority of first-place votes, so $c$ is the Coombs winner. But where $\mathbf{P}^{\prime}$ consists of each of the six linear orders over $\{a,b,c\}$, in $\mathbf{P}+\mathbf{P}^{\prime}$ no alternative receives a majority of first-place votes, and $c$ has the most last-place votes, so $c$ is eliminated in the first round (and then $a$ wins). Finally, for Mediancentre-Borda, start with the profile labeled “Irresolute OWM Failure” in the simulation accompanying Cervone et al. 2012 at https://omega.math.union.edu/research/2010-05-voting/; then adding one copy of each linear order causes the original winner to lose. However, clearly all scoring rules (Young 1975) and voting methods based on majority margins (see, e.g., Holliday and Pacuit 2023a) satisfy block preservation. Moreover, the methods cited above that violate block preservation also violate immunity to spoilers, so block preservation is not doing any special work by ruling out those methods.
We are now ready to state and prove our main result. [^12] Just before submitting the final version of this paper to the publisher, we learned from Chris Dong (personal communication) that our preprint of this paper (Holliday and Pacuit 2023c) inspired a related theorem in his recent co-authored preprint, Brandt et al. 2024. The theorem in Brandt et al. 2024 (Theorem 5) replaces May’s axioms and block preservation in Theorem 3.8 by an axiom of singleton negative involvement (see their Proposition 1): if $F(\mathbf{P})\neq\{a\}$, and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $a$ uniquely last, then $F(\mathbf{P}^{\prime})\neq\{a\}$. Their theorem also replaces near immunity to spoilers by Condorcet consistency, but note as in Proposition 3.10 below that Theorem 3.8 can be formulated in this way as well.
@@
Now suppose $\mathbf{P}$ is a three-alternative profile with no Condorcet winner. The margin graph of a profile $\mathbf{P}$ is the weighted directed graph with an edge from $a$ to $b$ if $a$ has a positive margin over $b$, weighted by that margin. It follows that the margin graph of $\mathbf{P}$ is of one of the following four forms, where the dashed arrows might not be present, due to a possibly zero margin, and the absence of an arrow indicates a zero margin:
@@
Note that we use ‘$a$’, ‘$b$’, and ‘$c$’ above as variables—so in the ascending case, we simply call the alternative with the largest loss ‘$a$’, etc. To see that the above four cases are exhaustive, assuming there is no Condorcet winner, observe that if the number of edges in the margin graph is zero, then we are in symmetric case; if the number of edges is one, then we are in the ascending case with $0=n=m$; if the number of edges is two, then we are in either the Condorcet loser case or the ascending case or descending case with $0=n<m$; and if the number of edges is three, then we are in either the ascending, descending, or symmetric case.
@@
Case 3: the margin graph of $\mathbf{P}$ is as in the Condorcet loser case. In this case, $Minimax(\mathbf{P})=\{b,c\}$, and the argument that $F(\mathbf{P})\subseteq\{b,c\}$ is the same as in the first paragraph of the ascending case, only $\mathrm{Margin}_{\mathbf{P}}(a,b)$ is negative, which is irrelevant to the argument.
@@
In Section 2, we assumed in the tradition of Arrow (1963) that the domain of a voting method is the set of profiles of strict weak orders. However, inspection of the proofs of Facts 3.1, 3.3, 3.4, and 3.7 and of Theorem 3.8 shows that for these results, we do not actually need to assume any properties of a voter’s binary relation $\mathbf{P}(i)$ except that it is a strict preference relation, i.e., it is asymmetric. [^13] When we say that $i$ “ranks $a$ uniquely first”, we simply mean that $(a,b)\in\mathbf{P}(i)$ for all $b\in X(\mathbf{P})\setminus\{a\}$ (and similarly for “ranks $a$ uniquely last” but with $(b,a)\in\mathbf{P}(i)$), not that $\mathbf{P}(i)$ forms a transitive ranking. In particular, all of these results also apply to the setting of linear ballots, where ties are not allowed, to the setting of truncated ballots, where some alternatives may be unranked (regarded as tied for last), and to the setting where instead of asking voters for a ranking, we simply ask for their pairwise choices (Do you prefer $a$ or $b$? Do you prefer $a$ or $c$? Do you prefer $b$ or $c$?), which allows for intransitivities in $\mathbf{P}(i)$. This latter point is especially important given the recent proposal of Foley and Maskin (2024) that in top-three general elections, we should ask voters to answer three pairwise comparison questions rather than to provide a ranking of the three alternatives. Theorem 3.8 continues to apply in this setting.
@@
Theorem 3.8 bears on a debate about how Minimax should be formulated. When an alternative $b$ has a positive margin over $a$, should we measure the size of $a$’s loss by the margin of $b$ over $a$, as we have, or rather by the support for $b$ over $a$, defined as the absolute number of voters who rank $b$ above $a$ (cf. Schulze 2011, p. 269)? If voters’ rankings are all strict linear orders, the choice does not matter, but in profiles allowing ties (even if only at the bottom of voters’ rankings), the two approaches are not equivalent. In fact, Minimax formulated in terms of support rather than margin violates positive involvement. Consider the following profile (where the number above each ranking indicates the number of voters with that ranking, and ‘$b,c$’ indicates a tie) with its margin graph (center) and support graph (right):
@@
Support-based Minimax says that $c$ has the “smallest loss”, so $c$ wins. But then as in the strictly ascending case of the proof of Theorem 3.8, adding $3$ voters with the ranking $cba$ makes $b$ the Condorcet winner and hence the support-based Minimax winner, violating positive involvement. We take this to be an argument in favor of the margin-based formulation of Minimax instead of the support-based formulation.
Finally, we note that there are indeed proper refinements of Minimax satisfying the axioms in Theorem 3.8. Since the issue of how to break Minimax ties is much less important than constraining a voting method to choose from among the Minimax winners, we defer discussion of such refinements to the Appendix. However, it is noteworthy that with the refinement defined in the Appendix, we satisfy not only weak positive responsiveness but even full positive responsiveness as in Remark 2.2. Yet Proposition A.2 in the Appendix shows that there is a tradeoff between satisfying positive responsiveness (rather than just weak positive responsiveness) and immunity to spoilers (rather than just near immunity to spoilers).
@@ 4 Independence of axioms
In this section, we turn to the question of the independence of the axioms in Theorem 3.8. First, however, we claim that we can drop May’s axiom of anonymity from our list of axioms. Anonymity is only used in the proof of Theorem 3.8 to help derive that the voting method refines majority voting in two-alternative profiles. But we can derive this result without anonymity as follows.
@@
Suppose $\mathbf{P}$ is a two-alternative profile with a unique majority winner. Hence there are $a,b\in X(\mathbf{P})$ such that $n$ voters rank $b$ above $a$, $n+k$ (for $k>0$) rank $a$ above $b$, and the remaining $\ell$ voters rank them as tied. Let $\mathbf{Q}=4\mathbf{P}$. If we can show that $F(\mathbf{Q})=\{a\}$, then by homogeneity, $F(\mathbf{P})=\{a\}$, in agreement with majority voting, as desired.
Let $\alpha=4n$, $\beta=4k$, and $\gamma=4\ell$. Since $k>0$, we have $\beta\geq 4$. In $\mathbf{Q}$, $\alpha$ voters rank $b$ above $a$, $\alpha+\beta$ rank $a$ above $b$, and the remaining $\gamma$ voters rank them as tied. Consider the profile $\mathbf{Q}_{0}$ consisting of just $\beta-2$ of the voters from $\mathbf{Q}$ who rank $a$ above $b$, still ranking $a$ above $b$, and all of the $\gamma$ voters from $\mathbf{Q}$ who ranked them as tied, still ranking them as tied. We claim that $F(\mathbf{Q}_{0})=\{a\}$. For if $b\in F(\mathbf{Q}_{0})$, then where $\mathbf{R}_{0}$ is obtained from $\mathbf{Q}_{0}$ by the $\beta-2$ voters who ranked $a$ above $b$ switching to rank $b$ above $a$, weak positive responsiveness applied $\beta-2$ times implies $F(\mathbf{R}_{0})=\{b\}$, while neutrality together with $b\in F(\mathbf{Q}_{0})$ implies $a\in F(\mathbf{R}_{0})$, a contradiction. Hence $F(\mathbf{Q}_{0})=\{a\}$. Now observe that there are exactly $\alpha+2$ voters from $\mathbf{Q}$ who ranked $a$ above $b$ who are not in $\mathbf{Q}_{0}$. Let $\mathbf{Q}_{1}$ be obtained from $\mathbf{Q}_{0}$ by adding all of those $\alpha+2$ voters, where $\alpha+1$ of them still rank $a$ above $b$ but one of them—say, voter $i_{1}$—now ranks $b$ above $a$, and adding all $\alpha$ of the original voters who ranked $b$ above $a$, still ranking $b$ above $a$. Then $\mathbf{Q}_{1}$ is obtained from $\mathbf{Q}_{0}$ by adding $\alpha+1$ pairs of the rankings $ab$ and $ba$, so by block preservation, $F(\mathbf{Q}_{0})=\{a\}$ implies $a\in F(\mathbf{Q}_{1})$. Finally, observe that $\mathbf{Q}$ can be obtained from $\mathbf{Q}_{1}$ by voter $i_{1}$ moving $a$ from last place to first place, so $a\in F(\mathbf{Q}_{1})$ implies $F(\mathbf{Q})=\{a\}$ by weak positive responsiveness, in agreement with majority voting. The relations between $\mathbf{Q}$, $\mathbf{Q}_{0}$, and $\mathbf{Q}_{1}$ are shown in Figure 1. ∎
@@
$\mathbf{Q}$ $h_{1}$dots $h_{\alpha}$ $i_{1}$dots $i_{\alpha+\beta}$ $j_{1}$dots $j_{\gamma}$ $b$dots $b$ $a$dots $a$ $a,b$dots $a,b$ $a$dots $a$ $b$dots $b$
$\mathbf{Q}_{0}$ $i_{\alpha+3}$dots $i_{\alpha+\beta}$ $j_{1}$dots $j_{\gamma}$ $a$dots $a$ $a,b$dots $a,b$ $b$dots $b$
$\mathbf{Q}_{1}$ $h_{1}$dots $h_{\alpha}$ $\boldsymbol{i_{1}}$ $i_{2}$dots $i_{\alpha+2}$ $i_{\alpha+3}$dots $i_{\alpha+\beta}$ $j_{1}$dots $j_{\gamma}$ $b$dots $b$ $\boldsymbol{b}$ $a$dots $a$ $a$dots $a$ $a,b$dots $a,b$ $a$dots $a$ $\boldsymbol{a}$ $b$dots $b$ $b$dots $b$ $\scriptstyle\alpha+1$ voters $\scriptstyle\alpha+1$ voters voters from $\scriptstyle\mathbf{Q}_{0}$
@@
We now show that each axiom in Theorem 3.8 is independent of the other axioms in the following sense.
@@
For each of the axioms $A$ in Theorem 3.8, there is a voting method $F_{A}$ on the domain of profiles with two or three alternatives such that $F_{A}$ satisfies all of the axioms in Theorem 3.8 except for $A$, which it violates; and if $A$ is any of the axioms other than anonymity, then $F_{A}$ does not refine Minimax.
@@
Where $A$ is anonymity, fix designated voters $i,j\in\mathcal{V}$ and define a voting method $F$ as follows: if $V(\mathbf{P})=\{i,j\}$ and $i$ and $j$ have reversed linear orders of the alternatives in $\mathbf{P}$, then $F(\mathbf{P})$ contains only $i$’s favorite alternative; otherwise $F(\mathbf{P})=Minimax(\mathbf{P})$. Clearly $F$ violates anonymity, but it is easy to check that it satisfies all the other axioms.
@@
Where $A$ is near immunity to spoilers, the Borda voting method clearly satisfies all the axioms except for near immunity to spoilers, which it violates in three-alternative profiles (see Holliday and Pacuit 2021a, Example 4.15); it is also easy to construct three-alternative profiles with $Borda(\mathbf{P})\not\subseteq Minimax(\mathbf{P})$. ∎
@@
In $\mathbf{P}$, there is no Condorcet winner, so $\mbox{{CP}}(\mathbf{P})=\{c\}$ since $c$ has the maximum plurality score of $5$. Then $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter with the ranking $cba$, but in $\mathbf{P^{\prime}}$, $b$ is the unique weak Condorcet winner, so $\mbox{{CP}}(\mathbf{P}^{\prime})=\{b\}$, violating positive involvement.
@@
Case 2: there is no weak Condorcet winner in $\mathbf{P}$. Then $a\in\mbox{{CP}}(\mathbf{P})$ implies that $a$ is among the alternatives with maximal plurality score in $\mathbf{P}$. Then clearly $a$ has the unique maximum plurality score in $\mathbf{P}^{\prime}$, and if moving from $\mathbf{P}$ to $\mathbf{P}^{\prime}$ creates a weak Condorcet winner, it must be $a$ alone, so in either case we have $\mbox{{CP}}(\mathbf{P}^{\prime})=\{a\}$.
@@
Observe that in $2\mathbf{P}$, the difference between the best and second best Minimax scores is 2, so $F(2\mathbf{P})=\{a\}$, violating homogeneity. Yet clearly $F$ satisfies anonymity, neutrality, block preservation (since adding a block of all linear orders does not change margins), and immunity to spoilers (since $F$ agrees with Minimax in two-alternative profiles and selects all of the Minimax winners in three-alternative profiles). It remains to show that $F$ satisfies weak positive responsiveness and positive involvement.
For positive involvement, for an arbitrary profile $\mathbf{P}$, suppose $x\in F(\mathbf{P})$ and that $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $x$ first. If $x\in Minimax(\mathbf{P})$, then $x\in Minimax(\mathbf{P}^{\prime})$ by positive involvement for Minimax, so $x\in F(\mathbf{P}^{\prime})$ by the definition of $F$. On the other hand, if conditions (I)-(III) hold and $x$ is the alternative with the second best Minimax score in $\mathbf{P}$, then since $x$ beats the Minimax winner $m$ head-to-head, $m$’s Minimax score is given by $x$’s margin over $m$ (since otherwise $m$ would have two losses, which is impossible for a Minimax winner in a three-alternative profile), so in $\mathbf{P}^{\prime}$ the Minimax score of $m$ increases by 1, while the Minimax score of $x$ decreases by $1$, so $x\in Minimax(\mathbf{P}^{\prime})\subseteq F(\mathbf{P}^{\prime})$.